# Suppressing Overestimation in Q-Learning through Adversarial Behaviors

### v1

@@ 4.2 Dummy Adversarial Q-Learning (DAQ)
${Q_{i}}({s_{t}},{a_{t}})\leftarrow{Q_{i}}({s_{t}},{a_{t}})+\alpha\left[r(s_{t},a_{t},s_{t+1})+{b_{i}}+\gamma{\max_{a^{\prime}\in{\mathcal{A}}}}{\min_{j\in\{1,\ldots,N\}}}{Q_{j}}(s_{t+1},a^{\prime})-{Q_{i}}(s_{t},a_{t})\right]$ (2)
@@
${Q_{i}}({s_{t}},{a_{t}})\leftarrow{Q_{i}}({s_{t}},{a_{t}})+\alpha\left[r(s_{t},a_{t},s_{t+1})+{b_{i}}+\gamma{\min_{j\in\{1,\ldots,N\}}}{\max_{a^{\prime}\in{\mathcal{A}}}}{Q_{j}}(s_{t+1},a^{\prime})-{Q_{i}}(s_{t},a_{t})\right]$ (3)
@@
The main difference from the maxmin and minmax Q-learning algorithms lies in the modified reward, $r(s_{t},a_{t},s_{t+1})+b_{i}$, where the constant $b_{i}$ depending on the Q-estimator index $i$ is added to the original reward when updating the target [^2] Unlike conventional reward shaping methods using shaping function $F(s,a,s^{\prime})$ which depends on state and action (Ng et al., 1999; Wiewiora et al., 2003), here we use a fixed constant.. The constant shift terms $b_{i}$ can be either positive or negative, and either random or deterministic, contingent upon the environment, functioning as design parameters. In subsequent sections, experimental results are presented to illustrate the impacts of different choices of $b_{i}$. Due to the constants, each estimator $Q_{i}$ learns a modified optimal Q-function
$Q_{i}^{*}(s,a):={\mathbb{E}}\left[{\left.{\sum\limits_{t=0}^{\infty}{{\gamma^{t}}(r({s_{t}},{a_{t}},{s_{t+1}})+{b_{i}})}}\right|{s_{0}}=s,{a_{0}}=s,{\pi^{*}}}\right]$
@@
${Q^{*}}(s,a):={\mathbb{E}}\left[{\left.{\sum\limits_{t=0}^{\infty}{{\gamma^{t}}r({s_{t}},{a_{t}},{s_{t+1}})}}\right|{s_{0}}=s,{a_{0}}=s,{\pi^{*}}}\right]$
with the constant bias $\sum_{t=0}^{\infty}{{\gamma^{t}}{{\mathbb{E}}[b_{i}]}}=\frac{{\mathbb{E}}[b_{i}]}{{1-\gamma}}$, i.e., $Q_{i}^{*}(s,a)-\frac{{\mathbb{E}}[b_{i}]}{{1-\gamma}}={Q^{*}}(s,a)$. Recalling that our goal is to find an optimal policy and finding the optimal Q-function aids this, adding a constant to the reward function does not affect the optimal policy which comes from a relative comparison among Q-function values of the given state, i.e., ${\pi^{*}}(s)=\arg{\max_{a\in{\mathcal{A}}}}{Q^{*}}(s,a)=\arg{\max_{a\in{\mathcal{A}}}}Q_{i}^{*}(s,a),i\in\{1,2,\ldots,N\}$.
A larger $N$ might lead to slower convergence but has the potential to significantly mitigate overestimation biases. The reduction in convergence speed can be counterbalanced by a synchronized update of the Q-iterate, a topic that will be elaborated upon in the ensuing subsection.
In summary, DAQ is capable of learning an optimal policy through the utilization of multiple Q-estimators without modifying the original objective. However, the adjustments made to the rewards and updates have the potential to regulate the overestimation biases arising from the max operator.
The fundamental intuition behind DAQ hinges on the viewpoint that it can be regarded as introducing a dummy player to the inherent Markov decision process, thereby establishing a two-player zero-sum Markov game (Shapley, 1953). Consequently, DAQ can be interpreted as Minimax Q-learning (Littman, 1994) for the Markov game, featuring a dummy player who, while not influencing the environment, can regulate the overestimation biases. These insights will be further detailed in the subsequent sections. Lastly, we foresee the extension of this approach to deep Q-learning and actor-critic algorithms as a viable means to augment their performances.
4.3 Discussion on Asynchronous versus Synchronous Updates
@@
4.4 Interpretation from the Two-Player Zero-Sum Game
DAQ can be interpreted as the Minimax Q-learning algorithm for a two-player zero-sum game (Lee, 2023; Shapley, 1953; Littman, 1994). The index $i$ of equation 2 and equation 3 can be interpreted as the action selected by the adversary, $r(s_{t},a_{t},s_{t+1})+b_{i}$ is the reward affected by both the user and adversary, and =${\mathcal{B}}=\{1,2,\dots,N\}$ is the action set of the adversary.
More specifically, consider the addition of a dummy player who, at each stage, executes an action $i\in{\mathcal{B}}$, incorporating the additional value $b_{i}$ into the reward $r(s_{t},a_{t},s_{t+1})$; that is, $r(s_{t},a_{t},s_{t+1})$ becomes $r({s_{t}},{a_{t}},{s_{t+1}})+{b_{i}}$. This player endeavors to minimize the return, the discounted summation of rewards, signifying its adversarial role, and thus reformulating the MDP into a two-player zero-sum game. The dummy player’s action does not alter the inherent transition probabilities of the MDP, preserving the user’s ability to attain its goal unimpeded. However, it can modify the overestimation biases during the learning updates, serving as the core concept of DAQ. The Minimax Q-learning algorithm, which is suitable for a two-player zero-sum game, can be adapted to this redefined game framework leading to
$Q({s_{t}},{a_{t}},i)\leftarrow Q({s_{t}},{a_{t}},i)+\alpha\left[{\underbrace{r({s_{t}},{a_{t}},i)}_{={r_{t}}+{b_{i}}}+\gamma{{\max}_{a\in{\mathcal{A}}}}{{\min}_{j\in\{1,\ldots,N\}}}Q({s_{t+1}},a,j)-Q({s_{t}},{a_{t}},i)}\right]$
@@ 4.4 Interpretation from the Two-Player Zero-Sum Game
$Q({s_{t}},{a_{t}},i)\leftarrow Q({s_{t}},{a_{t}},i)+\alpha\left[{\underbrace{r({s_{t}},{a_{t}},i)}_{={r_{t}}+{b_{i}}}+\gamma{{\min}_{j\in\{1,\ldots,N\}}}{{\max}_{a\in{\mathcal{A}}}}Q({s_{t+1}},a,j)-Q({s_{t}},{a_{t}},i)}\right]$
These algorithms are exactly identical to DAQ (minmax and maxmin versions) in this paper. Since the adversary tries to minimize the reward, she or he may cause underestimation biases, regulating the overestimation biases in the naïve Q-learning. We observe that the adversarial actions have no impact on the user’s goal as these actions are artificially incorporated into both the reward and the Q-updates, rendering them redundant. However, the incorporation of such dummy adversarial behavior can alter the overestimation biases.
@@
4.5 Finite-Time Convergence Analysis
Since DAQ can be interpreted as Minimax Q-learning for the two-player Markov game, existing analyses such as Lee (2023) and Littman & Szepesvári (1996) can be directly applied with minimal modifications. In this paper, we refer to the simple finite-time analysis presented in Lee (2023).
For simplicity of analysis, let us suppose that the step-size is a constant $\alpha\in(0,1)$, the initial iterate $Q$ satisfies $\left\|{Q}\right\|_{\infty}\leq 1$, $|r(s,a,s^{\prime})|\leq 1$ for all $(s,a,s^{\prime})\in{\cal S}\times{\cal A}\times{\cal S}$, and $\{(s_{k},a_{k},s_{k}^{\prime})\}_{k=0}^{\infty}$ are i.i.d. samples under the behavior policy $\beta$ for action $a\in{\cal A}$, where the behavior policy denotes the strategy by which the agent operates to collect experiences. Moreover, we assume that the state at each time is sampled from the stationary state distribution $p$, and in this case, the state-action distribution at each time is identically given by
$d(s,a)=p(s)\beta(a|s)\quad\forall(s,a,b)\in{\cal S}.$
Moreover, let us suppose that $d(s,a)>0$ holds for all $s\in{\cal S},a\in{\cal A}$. In the asynchronous version, the distribution $\mu$ of $i$ can be seen as the adversary’s behavior policy. We assume also that $\mu(i)>0,\forall i\in\{1,2,\ldots,N\}$. Moreover, let us define ${d_{\max}}:={\max_{(s,a,i)\in{\cal S}\times{\cal A}\times\{1,2,\ldots,N\}}}d(s,a)\mu(i)\in(0,1)$, ${d_{\min}}:={\min_{(s,a,i)\in{\cal S}\times{\cal A}\times\{1,2,\ldots,N\}}}d(s,a)\mu(i)\in(0,1)$, and $\rho:=1-\alpha d_{\min}(1-\gamma)$. Applying Theorem 4 from Lee (2023) directly yields the following theorem.
@@ 4.5 Finite-Time Convergence Analysis
${\mathbb{E}}[{\left\|{{Q_{i}}-Q_{i}^{*}}\right\|_{\infty}}]\leq$ $\frac{{27{d_{\max}}|{\cal S}\times{\cal A}|N{\alpha^{1/2}}}}{{{d_{{{\min}^{3/2}}}}{{(1-\gamma)}^{5/2}}}}+\frac{{6|{\cal S}\times{\cal A}{|^{3/2}}{N^{3/2}}}}{1-\gamma}{\rho^{t}}$
$+\frac{{24\gamma{d_{\max}}|{\cal S}\times{\cal A}{|^{2/3}}{N^{2/3}}}}{{1-\gamma}}\frac{3}{{{d_{\min}}(1-\gamma)}}{\rho^{t/2-1}},$
@@
Note that in the bound detailed in Theorem 1, both the second and third terms diminish exponentially as the iteration step $t$ approaches infinity. The first term stands as a constant bias term, originating from the use of the constant step-size $\alpha$, but this term disappears as $\alpha\to 0$. This suggests that by employing a sufficiently small step-size, the constant bias can be effectively controlled. Along similar lines, the finite-time error bound for the synchronous version can be derived as in Theorem 1 with $d_{\max}$ and $d_{\min}$ replaced with ${d_{\max}}:={\max_{(s,a,i)\in{\cal S}\times{\cal A}}}d(s,a)\in(0,1)$ and ${d_{\min}}:={\min_{(s,a,i)\in{\cal S}\times{\cal A}}}d(s,a)\in(0,1)$, respectively. Note also that in this case, the error bound in Theorem 1 becomes smaller because $d_{\min}$ tends to increase and $d_{\max}$ tends to decrease in the synchronous case.
5 Experiments
@@
In this section, the efficacy of DAQ is demonstrated across several benchmark environments, and a comparative analysis is provided between DAQ and existing algorithms, including standard Q-learning, Double Q-learning, maxmin Q-learning, and minmax Q-learning. An $\epsilon$-greedy behavior based on $\sum Q_{i}$ and tabular action-values initialized with 0 are used. Any ties in $\epsilon$-greedy action selection are broken randomly.
5.1 MDP Environments
@@
Grid World in Hasselt (2010) is considered, which is illustrated in Figure 0(a). In this environment, the agent can take one of four actions at each state: move up, down, left, or right. The episode starts at $\mathsf{S}$, and the goal state is $\mathsf{G}$. For each transition to non-terminal states, the agent receives a random reward $r=-12$ or $r=+10$ with equal probabilities. When transiting to the goal state, every action incurs $r=+5$, and the episode is terminated. If the agent follows an optimal policy, the episode is terminated after five successive actions, and hence, the expected reward per step is $+0.2$. We will also consider another reward function adapted from Wang et al. (2020). For each transition, the agent receives a fixed reward $r=-1$. When transiting to the goal state, every action yields $r=-35$ or $r=+45$ with equal probabilities, and the episode ends. Under an optimal policy, the expected average reward per step is also $+0.2$.
@@
Sutton’s Example shown Figure 6.5 of Sutton & Barto (2018) is illustrated in Figure 0(b). There exist two non-terminal states $\mathsf{A}$ and $\mathsf{B}$ at which two actions, $\mathsf{left}$ and $\mathsf{right}$, are admitted. There are two terminal states denoted with grey boxes and $\mathsf{A}$ is the initial state. Executing the action $\mathsf{right}$ leads to the termination of an episode with no rewards, while the action $\mathsf{left}$ leads to state $\mathsf{B}$ with no reward. There exist many actions at state $\mathsf{B}$, which will all immediately terminate the episode with a reward drawn from a normal distribution $\mathcal{N}(-0.1,1)$. In our experiments, we will consider the two cases $r\sim\mathcal{N}(\mu,1)$ with $\mu=-0.1$ and $\mu=+0.1$.
@@
Weng’s Example from Weng et al. (2020) is adapted, and it is illustrated in Figure 0(c). There exist $M+1$ states labeled as $\{\sf{0},\dots,\sf{M}\}$ with two actions, $\mathsf{left}$ and $\mathsf{right}$, at each state, where $\mathsf{0}$ is the initial state. Choosing action $\mathsf{right}$ at the initial state leads to the terminal state, while choosing $\mathsf{left}$ leads to random transitions to one of the other $M$ states. At the initial state, both actions result in the reward $r=0$. At any $s\in\{\mathsf{1},\mathsf{2},\ldots,\mathsf{M}\}$, it transits to $\mathsf{0}$ with the action $\mathsf{right}$. On the other hand, the episode is terminated with the action $\mathsf{left}$. Both actions incur a reward drawn from a normal distribution $\mathcal{N}(-0.1,1)$.
(a) Grid World
(b) Sutton’s Example
(c) Weng’s Example
Figure 1: MDP Environments
5.2 Empirical Results

### v3

@@ IV-B Dummy Adversarial Q-Learning (DAQ)
${Q_{i}}(s_{t},a_{t})\leftarrow{Q_{i}}(s_{t},a_{t})+\alpha\biggl{[}r(s_{t},a_{t},s_{t+1})+b_{i}+\\ \gamma{\max_{a^{\prime}\in\mathcal{A}}}{\min_{j\in\{1,\ldots,N\}}}{Q_{j}}(s_{t+1},a^{\prime})-{Q_{i}}(s_{t},a_{t})\biggr{]}$ (2)
@@
${Q_{i}}(s_{t},a_{t})\leftarrow{Q_{i}}(s_{t},a_{t})+\alpha\biggl{[}r(s_{t},a_{t},s_{t+1})+b_{i}+\\ \gamma{\min_{j\in\{1,\ldots,N\}}}{\max_{a^{\prime}\in\mathcal{A}}}{Q_{j}}(s_{t+1},a^{\prime})-{Q_{i}}(s_{t},a_{t})\biggr{]}$ (3)
@@
The main difference from the maxmin and minmax Q-learning algorithms lies in the modified reward, $r(s_{t},a_{t},s_{t+1})+b_{i}$, where the constant $b_{i}$ depending on the Q-estimator index $i$ is added to the original reward when updating the target [^2] Unlike conventional reward shaping methods using shaping function $F(s,a,s^{\prime})$ which depends on state and action [11, 22], here we use a fixed constant.. The constant shift terms $b_{i}$ can be either positive or negative, and either random or deterministic, contingent upon the environment, functioning as design parameters. In subsequent sections, experimental results are presented to illustrate the impacts of different choices of $b_{i}$. Due to the constants, each estimator $Q_{i}$ learns a modified optimal Q-function
$Q_{i}^{*}(s,a):=\mathbb{E}\Biggl{[}{\sum\limits_{t=0}^{\infty}{{\gamma^{t}}(r(s_{t},a_{t},s_{t+1})+b_{i})}}\Bigg{|}\\ {s_{0}}=s,{a_{0}}=s,{\pi^{*}}\Biggr{]}$
@@
${Q^{*}}(s,a):=\mathbb{E}\Biggl{[}{\sum\limits_{t=0}^{\infty}{{\gamma^{t}}r(s_{t},a_{t},s_{t+1})}}\Bigg{|}\\ {s_{0}}=s,{a_{0}}=s,{\pi^{*}}\Biggr{]}$
with the constant bias $\sum_{t=0}^{\infty}{{\gamma^{t}}{{\mathbb{E}}[b_{i}]}}=\frac{{\mathbb{E}}[b_{i}]}{{1-\gamma}}$, i.e., $Q_{i}^{*}(s,a)-\frac{{\mathbb{E}}[b_{i}]}{{1-\gamma}}={Q^{*}}(s,a)$. Recalling that our goal is to find an optimal policy and finding the optimal Q-function aids this, adding a constant to the reward function does not affect the optimal policy which comes from a relative comparison among Q-function values of the given state, i.e., ${\pi^{*}}(s)=\operatorname*{arg\!max}_{a\in\mathcal{A}}{Q^{*}}(s,a)=\operatorname*{arg\!max}_{a\in\mathcal{A}}Q_{i}^{*}(s,a),i\in\{1,2,\ldots,N\}$.
A larger $N$ might lead to slower convergence but has the potential to significantly mitigate overestimation biases. The reduction in convergence speed can be counterbalanced by a synchronized update of the Q-iterate, a topic that will be elaborated upon in the following subsections.
In summary, DAQ is capable of learning an optimal policy through the use of multiple Q-estimators without modifying the original objective. However, the adjustments made to the rewards and updates have the potential to regulate the overestimation biases arising from the max operator.
The fundamental intuition behind DAQ hinges on the viewpoint that it can be regarded as introducing a dummy player to the inherent Markov decision process, thereby establishing a two-player zero-sum Markov game [15]. Consequently, DAQ can be interpreted as minimax Q-learning [9] for the Markov game, featuring a dummy player who, while not influencing the environment, can regulate the overestimation biases. These insights will be further detailed in the subsequent sections. Lastly, we foresee the extension of this approach to deep Q-learning and actor-critic algorithms as viable methods to improve their performances.
IV-C Discussion on Asynchronous versus Synchronous Updates
@@
IV-D Interpretation from the Two-Player Zero-Sum Game
DAQ can be interpreted as the minimax Q-learning algorithm for a two-player zero-sum game [6, 15, 9]. The index $i$ of (2) and (3) can be interpreted as the action selected by the adversary, $r(s_{t},a_{t},s_{t+1})+b_{i}$ is the reward affected by both the user and adversary, and =$\mathcal{B}=\{1,2,\dots,N\}$ is the action set of the adversary.
More specifically, let us consider the addition of a dummy player who, at each stage, takes an action $i\in\mathcal{B}$, incorporating the additional value $b_{i}$ into the reward $r(s_{t},a_{t},s_{t+1})$; that is, $r(s_{t},a_{t},s_{t+1})$ becomes $r({s_{t}},{a_{t}},{s_{t+1}})+{b_{i}}$. This player wants to minimize the return, the discounted summation of rewards, which signifies its adversarial role, and thus concentually reformulates the MDP into a two-player zero-sum game. The dummy player’s action does not alter the inherent transition probabilities of the MDP, and preserves the user’s ability to attain its goal unimpeded. However, it can modify the overestimation biases during the learning updates, and serves as the core concept of DAQ. The minimax Q-learning algorithm, which is suitable for a two-player zero-sum game, can be adapted to this redefined game framework, leading to
$Q(s_{t},a_{t},i)\leftarrow Q(s_{t},a_{t},i)+\alpha\Biggl{[}\underbrace{r(s_{t},a_{t},i)}_{=r_{t}+b_{i}}+\\ \gamma{{\max}_{a\in\mathcal{A}}}{{\min}_{j\in\{1,\ldots,N\}}}Q({s_{t+1}},a,j)-Q(s_{t},a_{t},i)\Biggr{]}$
@@ IV-D Interpretation from the Two-Player Zero-Sum Game
$Q(s_{t},a_{t},i)\leftarrow Q(s_{t},a_{t},i)+\alpha\Biggl{[}\underbrace{r(s_{t},a_{t},i)}_{=r_{t}+b_{i}}+\\ \gamma{{\min}_{j\in\{1,\ldots,N\}}}{{\max}_{a\in\mathcal{A}}}Q({s_{t+1}},a,j)-Q(s_{t},a_{t},i)\Biggr{]}$
These algorithms are exactly identical to DAQ (minmax and maxmin versions) in this paper. Since the adversary tries to minimize the reward, she or he may cause underestimation biases, regulating the overestimation biases in the Q-learning. We observe that the adversarial actions have no impact on the user’s goal as these actions are artificially incorporated into both the reward and the Q-updates, which means that they are redundant. However, the incorporation of such dummy adversarial behavior can alter the overestimation biases.
@@
IV-E Finite-Time Convergence Analysis
Since DAQ can be interpreted as minimax Q-learning for the two-player Markov game, existing analyses such as [6] and [10] can be directly applied with minimal modifications. In this paper, we refer to the simple finite-time analysis presented in [6].
For simplicity of analysis, let us suppose that the step-size is a constant $\alpha\in(0,1)$, the initial iterate $Q$ satisfies $\left\|{Q}\right\|_{\infty}\leq 1$, $|r(s,a,s^{\prime})|\leq 1$ for all $(s,a,s^{\prime})\in{\mathcal{S}}\times{\mathcal{A}}\times{\mathcal{S}}$, and $\{(s_{k},a_{k},s_{k}^{\prime})\}_{k=0}^{\infty}$ are i.i.d. samples under the behavior policy $\beta$ for action $a\in{\mathcal{A}}$, where the behavior policy denotes the strategy by which the agent operates to collect experiences. Moreover, we assume that the state at each time is sampled from the stationary state distribution $p$, and in this case, the state-action distribution at each time is identically given by
$d(s,a)=p(s)\beta(a|s)\quad\forall(s,a,b)\in{\mathcal{S}}.$
Moreover, let us suppose that $d(s,a)>0$ holds for all $s\in{\mathcal{S}},a\in{\mathcal{A}}$. In the asynchronous version, the distribution $\mu$ of $i$ can be seen as the adversary’s behavior policy. We assume also that $\mu(i)>0,\forall i\in\{1,2,\ldots,N\}$. Moreover, let us define ${d_{\max}}:={\max_{(s,a,i)\in{\mathcal{S}}\times{\mathcal{A}}\times\{1,2,\ldots,N\}}}d(s,a)\mu(i)\in(0,1)$, ${d_{\min}}:={\min_{(s,a,i)\in{\mathcal{S}}\times{\mathcal{A}}\times\{1,2,\ldots,N\}}}d(s,a)\mu(i)\in(0,1)$, and $\rho:=1-\alpha d_{\min}(1-\gamma)$. Applying Theorem 4 from [6] directly yields the following theorem.
@@ IV-E Finite-Time Convergence Analysis
$\mathbb{E}[\left\|Q_{i}-Q_{i}^{*}\right\|_{\infty}]\leq$ $\frac{27{d_{\max}}|\mathcal{S}\times\mathcal{A}|N{\alpha^{1/2}}}{{d_{{\min}^{3/2}}}{{(1-\gamma)}^{5/2}}}$
$+\frac{{6|\mathcal{S}\times{\mathcal{A}}{|^{3/2}}{N^{3/2}}}}{1-\gamma}{\rho^{t}}$
$+\frac{{24\gamma{d_{\max}}|\mathcal{S}\times\mathcal{A}{|^{2/3}}{N^{2/3}}}}{{1-\gamma}}$
×$\times\frac{3}{{{d_{\min}}(1-\gamma)}}{\rho^{t/2-1}},$
@@
Note that in the bound detailed in Theorem 1, both the second and third terms diminish exponentially as the iteration step $t$ approaches infinity. The first term can be seen as a constant bias term, which comes from the use of the constant step-size $\alpha$, but this term disappears as $\alpha\to 0$. This suggests that by employing a sufficiently small step-size, the constant bias can be effectively controlled. Along similar lines, the finite-time error bound for the synchronous version can be derived as in Theorem 1 with $d_{\max}$ and $d_{\min}$ replaced with ${d_{\max}}:={\max_{(s,a,i)\in{\mathcal{S}}\times{\mathcal{A}}}}d(s,a)\in(0,1)$ and ${d_{\min}}:={\min_{(s,a,i)\in{\mathcal{S}}\times{\mathcal{A}}}}d(s,a)\in(0,1)$, respectively. Note also that in this case, the error bound in Theorem 1 becomes smaller because $d_{\min}$ tends to increase and $d_{\max}$ tends to decrease in the synchronous case.
V Experiments
@@
In this section, the efficacy of DAQ is demonstrated across several benchmark environments, and a comparative analysis is provided between DAQ and existing algorithms, including standard Q-learning, double Q-learning, maxmin Q-learning, and minmax Q-learning. An $\epsilon$-greedy behavior based on $\sum Q_{i}$ and tabular action-values initialized with 0 are used. Any ties in $\epsilon$-greedy action selection are broken randomly. This random tie-breaking ensures that all actions with the same highest predicted value have an equal chance of being selected, promoting exploration and maintaining a balance between exploration and exploitation.
V-A MDP Environments
@@
Grid World of [3] is considered, which is illustrated in Fig. 1. In this environment, the agent can take one of four actions at each state: move up, down, left, or right. The episode starts at $\mathsf{S}$, and the goal state is $\mathsf{G}$. For each transition to non-terminal states, the agent receives a random reward $r=-12$ or $r=+10$ with equal probabilities. When transiting to the goal state, every action incurs $r=+5$, and the episode is terminated. If the agent follows an optimal policy, the episode is terminated after five successive actions, and hence, the expected reward per step is $+0.2$. We will also consider another reward function adapted from [19]. For each transition, the agent receives a fixed reward $r=-1$. When transiting to the goal state, every action yields $r=-35$ or $r=+45$ with equal probabilities, and the episode ends. Under an optimal policy, the expected average reward per step is also $+0.2$. Hereafter, we call these reward functions as $r^{H}_{t}$ and $r^{W}_{t}$, respectively.
$r^{H}_{t}=\begin{cases}+5&:\textit{at terminating step}\\ -12\textit{ or }+10&:\textit{otherwise}\end{cases}$
$r^{W}_{t}=\begin{cases}-35\textit{ or }+45&:\textit{at terminating step}\\ -1&:\textit{otherwise}\end{cases}$
Figure 1: MDP Environments - (a) Grid World
@@
Sutton’s MDP for an example of maximization bias in [16] is a well-known example where maximization bias degrades the performance of Q-learning while double Q-learning is not affected. There exist two non-terminal states $\mathsf{A}$ and $\mathsf{B}$ at which two actions, $\mathsf{left}$ and $\mathsf{right}$, are admitted as illustrated in Fig. 2. There also are two terminal states denoted with grey boxes and $\mathsf{A}$ is the initial state. Executing the action $\mathsf{right}$ leads to the termination of an episode with no rewards, while the action $\mathsf{left}$ leads to state $\mathsf{B}$ with no reward. There exist many actions at state $\mathsf{B}$, which will all immediately terminate the episode with a reward drawn from a normal distribution $\mathcal{N}(-0.1,1)$. In our experiments, we will consider the two cases $r\sim\mathcal{N}(\mu,1)$ with $\mu=-0.1$ and $\mu=+0.1$.
Figure 2: MDP Environments - (b) Sutton’s MDP
@@
Weng’s MDP for another example of maximization bias of Q-learning from [21] is adapted, and it is illustrated in Fig. 3. There exist $M+1$ states labeled as $\{\sf{0},\dots,\sf{M}\}$ with two actions, $\mathsf{left}$ and $\mathsf{right}$, at each state, where $\mathsf{0}$ is the initial state. Choosing action $\mathsf{right}$ at the initial state leads to the terminal state, while choosing $\mathsf{left}$ leads to random transitions to one of the other $M$ states. At the initial state, both actions result in the reward $r=0$. At any $s\in\{\mathsf{1},\mathsf{2},\ldots,\mathsf{M}\}$, it transits to $\mathsf{0}$ with the action $\mathsf{right}$. On the other hand, the episode is terminated with the action $\mathsf{left}$. Both actions incur a reward drawn from a normal distribution $\mathcal{N}(-0.1,1)$.
Figure 3: MDP Environments - (c) Weng’s MDP
V-B Empirical Results
Here, we demonstrate the superiority of DAQs through various results. Results are averaged over 10K runs.